OpenStudy (anonymous):

What are the solutions of the equation? z^2 – 10z + 24 = 0 –4, –6 4, –6 4, 6 –4, 6

OpenStudy (phi):

There are 2 ways to do this: 1) factor the quadratic or 2) test the answers. only 1 pair of the 4 will work.

OpenStudy (anonymous):

My answer is C.

OpenStudy (phi):

Do you want me to check if you put either 4 or 6 into the equation in place of z, you will get 0 ?

OpenStudy (anonymous):

That would be helpful i suppose. Been doing questions like these almost all day.

OpenStudy (anonymous):

I can never seem to fully understand it.

OpenStudy (phi):

Here is replacing z with 4 z^2 – 10z + 24 = 0 4^2 - 10(4) + 24 = ? 0 16 - 40 + 24 = ? 0 yes it works for 4 show me your work for replace z with 6

OpenStudy (anonymous):

Its all on paper, and i really don't want to transfer all of what i wrote onto this, it would be quite time consuming. Plus i am in a hurry because i have to go out with my family for the rest of the day.

OpenStudy (anonymous):

Can you simply tell me if i am incorrect? I'll continue the problem myself after.